OpenStudy (hexagon001):

find the residue of the following function at the point 0 f(z)=(1+z^2)/z can you tell me how its worked out, please

OpenStudy (anonymous):

at z=0 function has a simple pole, so residue will be 1

OpenStudy (hexagon001):

can you show me what the laurent series will be of this, please?

OpenStudy (anonymous):

f(z) can be written like: f(z)=1/z +z and this is it's Laurent series

OpenStudy (anonymous):

Ok?

OpenStudy (hexagon001):

thanks.i understand. my book essentially says the same thing but it shows the following.. (1+z^2)/z =..+0/z^3 +0/z^2 + 1/z +0 +z +0z^2... I am not understanding why they are including all the other bits inbetween and how and why they come about... could oyu explain please

OpenStudy (anonymous):

I think this is by definition: \[a_n=\frac{1}{2\pi i}\int\limits_\gamma\frac{f(z)}{(z-c)^{n+1}}\]

OpenStudy (hexagon001):

so for the same function in the question..for alpha=0..how would i work out \[\alpha _{-1}/(z-\alpha\])

OpenStudy (anonymous):

but your function is the case where \(\alpha=0 \)

OpenStudy (hexagon001):

ok ignoring alpha =0

OpenStudy (anonymous):

Or I am not understanding right your question....

OpenStudy (hexagon001):

how would i work that α−1/(z−α )

OpenStudy (anonymous):

what you need?

OpenStudy (hexagon001):

according to my book.. the residue of an analytic function f at a singularity α of f is the coeffiecient \[\alpha _{-1}\] of \[(z-\alpha)^{-1}\] in the laurent series for f : f(z)=\[...+\alpha _{-2}/(z-\alpha) +\alpha_{-1}/(z-\alpha) +\alpha_{0} +\alpha_{1}(z-\alpha)+...\], we can determine Res(f,alpha) by inspecting the series.. now my question specifically is how would i work out the coeffecient \[\alpha _{-1}\]

OpenStudy (anonymous):

By finding the Laurent seies, :)

OpenStudy (hexagon001):

and that is by using the definition you stated a few posts up right?

OpenStudy (anonymous):

there are also two formulas: http://en.wikipedia.org/wiki/Residue_(complex_analysis)#Limit_formula_for_higher_order_poles

OpenStudy (anonymous):

for simple and higher order poles. Or ya, by using formula I gave you above. But it can be tough to calculate the actual integral. Normaly finding Laurent series is easyer

OpenStudy (hexagon001):

aaargghh.. this is really chewing my brain out... I think i get it...i'll have to attempt a few questions and maybe do some more reading to get my head around it. Thanks Myko..you are a star for helping. wish i could give you more medals..

OpenStudy (anonymous):

:). Good luck. Complex analysis is tough at the beginning